# Editing the Mind of Giants: An In-Depth Exploration of Pitfalls of Knowledge Editing in Large Language Models

### v1

@@ Abstract
Knowledge editing is a rising technique for efficiently updating factual knowledge in Large Language Models (LLMs) with minimal alteration of parameters. However, recent studies have identified concerning side effects, such as knowledge distortion and the deterioration of general abilities, that have emerged after editing. This survey presents a comprehensive study of these side effects, providing a unified view of the challenges associated with knowledge editing in LLMs. We discuss related works and summarize potential research directions to overcome these limitations. Our work highlights the limitations of current knowledge editing methods, emphasizing the need for deeper understanding of inner knowledge structures of LLMs and improved knowledge editing methods. To foster future research, we have released the complementary materials such as paper collection publicly [^1] https://github.com/MiuLab/EditLLM-Survey. **footnotetext: Equal contribution.
@@
Cheng-Hsun Hsueh Paul Kuo-Ming Huang Tzu-Han Lin Che-Wei Liao Hung-Chieh Fang Chao-Wei Huang Yun-Nung Chen National Taiwan University, Taipei, Taiwan {r12922059, b08902072, b08902126, r09922a25}@csie.ntu.edu.tw {b09902106, f07922069}@csie.ntu.edu.tw y.v.chen@ieee.org
@@ 1 Introduction
Recent advancements in large language models (LLMs) have significantly improved Natural Language Processing (NLP; Brown et al., 2020). LLMs can now understand and generate language at a human-like level, establishing them as milestones in generative AI. This proficiency is due to their ability to extract knowledge from extensive text corpora. However, the mechanisms of knowledge storage in LLMs are not well understood, leading to potential issues in real-world applications. For instance, deploying LLMs as chatbots raises concerns about the reliability of their generated content due to the unclear mechanisms of knowledge storage and potential inaccuracies.
Figure 1: An overview of pitfalls in current knowledge editing methods. The subsequent sections dive into three key challenges: generalization issues (Section 3.1), localization issues (Section 3.2), and deterioration of general LLM abilities (Section 3.3).
@@
To address these issues, researchers have explored various methods. Traditional approaches like fine-tuning, continual learning, and retraining are either computationally expensive or detrimental to the LLMs’ overall performance. Recently, knowledge editing has emerged as a promising alternative, offering minimal computational costs and fewer alterations (Cao et al., 2021; Dai et al., 2022; Meng et al., 2022, 2023; Dong et al., 2022; Mitchell et al., 2022a, b; Hartvigsen et al., 2023; Huang et al., 2023; Yu et al., 2024; Zheng et al., 2023; Li et al., 2024a; Tan et al., 2024; Gupta et al., 2024b; Wang et al., 2024). This approach allows for precise adjustments to LLMs, enhancing their practicality and reliability in real-world applications.
@@
Knowledge editing can be divided into two main categories: parameter-modifying and parameter-preserving. Both aim to refine LLM knowledge efficiently while avoiding the drawbacks of previous tuning methods (Yao et al., 2023a). Parameter-modifying methods, including meta-learning (Cao et al., 2021; Mitchell et al., 2022a; Tan et al., 2024) and locate-then-edit techniques (Dai et al., 2022; Meng et al., 2022, 2023; Li et al., 2024a; Gupta et al., 2024b), strive to update a targeted set of model parameters while making effective edits. Parameter-preserving methods introduce external components, like knowledge bases (Mitchell et al., 2022b; Yu et al., 2024) or extra model parameters (Dong et al., 2022; Huang et al., 2023; Hartvigsen et al., 2023; Yu et al., 2024) to maintain the integrity of pre-trained LLMs while updating their knowledge.
Despite the success of knowledge editing, challenges remain. Knowledge editing can have unintended side effects, potentially damaging the general abilities and intrinsic structures of LLMs. Previous research has mainly focused on performance improvements and innovations within knowledge editing methods, with limited attention to potential drawbacks. Consequently, this survey aims to provide a holistic view of current issues in the knowledge editing paradigm and encourage further investigations into the pitfalls and intrinsic knowledge structures of LLMs. A brief overview of the discussed pitfalls is shown in Figure  1.
@@
This survey is organized as follows: Section 2 introduces knowledge editing and its methods. Section 3 discusses current challenges, benchmarks, and methodologies. In Section 4, we present experimental results evaluating different editing methods. Finally, Section 5 explores related studies and future research directions. We summarize our contributions as follows:
@@
    We systematically organized previous research and conducted experiments to benchmark the side effects of knowledge editing, providing a unified perspective on this issue.
@@
    We discussed related studies and potential research directions to address existing challenges, encouraging further exploration and understanding in this field.
Figure 2: Illustration of properties that knowledge editing methods should satisfy. An ideal knowledge editing method should be reliable, generalizable to relevant queries, and should not alter the outputs of irrelevant queries.
@@
Category Strategy Method
MEND (Mitchell et al., 2022a)
MALMEN (Tan et al., 2024)
Locating and editing Knowledge Neuron (Dai et al., 2022)
ROME (Meng et al., 2022)
MEMIT (Meng et al., 2023)
PMET (Li et al., 2024a)
EMMET (Gupta et al., 2024b)
T-Patcher (Huang et al., 2023)
GRACE (Hartvigsen et al., 2023)
MELO (Yu et al., 2024)
External memory SERAC (Mitchell et al., 2022b)
MeLLo (Zhong et al., 2023)
In-context learning IKE (Zheng et al., 2023)
Decoding DeepEdit (Wang et al., 2024)
Table 1: Overview of current knowledge editing methods. The methods are categorized into two major families, namely parameter-modifying and parameter-preserving methods, each containing several strategies. Methods marked with $\dagger$ have the ability to process sequential edits.
@@ 2.1 Problem Definition
Knowledge editing for LLMs entails modifying the output of LLMs in response to specific edit queries, with the aim of minimizing alterations to their original behavior (Yao et al., 2023a; Mazzia et al., 2023; Zhang et al., 2024a). In this section, we follow the notation from Mazzia et al. (2023).
@@
The editing process should not affect instances unrelated to the edit queries. The locality set of an edit query $(x_{e},y_{e})$ can be defined as $L(x_{e})=\{(x_{loc},y_{loc})\in\mathbb{X}\times\mathbb{Y}\ \mathrm{s.t}\ x_{loc}\notin N(x_{e},y_{e})\land f_{\theta_{0}}(x_{loc})=y_{loc}\}$. The locality, also known as specificity, of a editing method is measured by calculating the level of invariance of model output before and after the edits, which can be calculated as follows:
@@ 2.2 Current Methods
In this section, we introduce the current knowledge editing methods. The methods are categorized into parameter-modifying (Section 2.2.1) and parameter-preserving (Section 2.2.2) editing methods, each containing several strategies. An overview and illustration of current methods are included in Table 1 and Figure 3, respectively.
@@ 2.2.1 Parameter-Modifying
This category of methods, including meta-learning methods and locate-and-edit strategies, update LLMs’ knowledge by modifying their parameters.
@@
Meta-learning methods train a hyper-network to predict the update of network parameters. For instance, KnowledgeEditor (Cao et al., 2021) trains a deep network to predict weight updates. MEND (Mitchell et al., 2022a) decomposes the gradient matrix into two rank-one matrices and utilized a hyper-network to update these matrices, thereby accelerating the editing process. Built upon MEND, MALMEN (Tan et al., 2024) refines the process by formulating the aggregation of parameter shifts into a least-squares problem, further improving the scalability of meta-learning methods.
@@
Locate-and-edit methods identify specific knowledge locations in LLMs for consequent editing. KN (Dai et al., 2022) utilizes the proposed knowledge attribution method to pinpoint neurons expressing relational facts, allowing efficient updates or erasures without fine-tuning. ROME (Meng et al., 2022) proposes the causal tracing method for identifying neuron activations associating with specific knowledge. The authors demonstrate the significance of middle-layer feed-forward networks (FFNs) in factual predictions when processing the subject’s last token. Built upon the hypothesis that the FFN modules in a transformer layer can be viewed as key-value memories (Geva et al., 2021), ROME (Meng et al., 2022) injects new knowledge into the key-value memories by deriving the closed form solution from the least-squares problem. MEMIT (Meng et al., 2023) scales up ROME (Meng et al., 2022) by editing a set of MLPs from consecutive middle-layers via solving a normal equation. PMET (Li et al., 2024a) proposes to update multi-head self-attention (MHSA) modules in addition to FFNs. EMMET (Gupta et al., 2024b) on the other hand, integrates the objectives of ROME and MEMIT into a unified preservation-memorization objective, facilitating batch-editing capabilities for both methodologies.
@@ 2.2.2 Parameter-Preserving
Parameter-preserving methods alter LLM output by adding new parameters, integrating external memory, or using strategies like in-context learning and devised decoding, without changing pre-trained LLM.
@@
Some methods utilize additional parameters, such as adding new neurons or employing parameter-efficient techniques. CaliNET (Dong et al., 2022) extends the FFN modules with calibration memory slots to adjust the predicted token distribution. T-Patcher (Huang et al., 2023) adds neurons in the FFN’s last layer to rectify classification errors and incorrectly generated tokens, activating only in response to associated mistakes. GRACE (Hartvigsen et al., 2023) wraps a selected layer with an Adaptor that includes a codebook and deferral mechanism, learning to decode desired outputs while caching input error embeddings. The GRACE layer stores the edits and could be updated continuously over long deployments. MELO (Yu et al., 2024) utilizes DyLoRA (Valipour et al., 2023) modules to learn edits, indexing them in an inner vector database to dynamically activate corresponding LoRA blocks during inference.
@@ 3 Challenges of Knowledge Editing
While knowledge editing methods have been extensively researched, there’s a lack of comprehensive study on related challenges. In this section, we discuss the pitfalls of knowledge editing from three perspectives: inability to logically inference and robustly generalize (Section 3.1), unintended alteration of non-target knowledge (Section 3.2), and deterioration of general LLM abilities (Section 3.3).
@@
Portability and Generalization RippleEdits (Cohen et al., 2023) Logical Generalization
Compositionality I
Compositionality II
ConflictEdit (Li et al., 2024b) Conflict Score
Conflict Magnitude
Success Score
MQuAKE (Zhong et al., 2023) Edit-wise Success Rate
Instance-wise Accuracy
Multi-hop Accuracy
@@
ZsRE + CounterFact Yao et al. (2023b) Subject-Replace
Reverse-Relation
One-Hop
Locality RippleEdits (Cohen et al., 2023) Subject Aliasing
Preservation
Relation Specificity
RoundEdit (Li et al., 2024b) Success Score
Distortion ($\downarrow$)
Ignore Rate ($\downarrow$)
Failure Rate ($\downarrow$)
Tied Fact Damage ($\downarrow$)
CounterFact Yao et al. (2023b) Other-Attribution
Distract-Neighbor
Other-Task
CounterFact Meng et al. (2022) Locality
Neighborhood Score
Neighborhood Magnitude
@@
Table 2: Performance benchmarks and evaluation metrics addressing generalization/portability and locality issues in knowledge editing methods. Unless specifically indicated by a downward arrow, higher values signify better performance in those evaluation metrics. CounterFact benchmark is proposed by  Meng et al. (2022), and CounterFact with mark is modified to further examine the proposed metrics.
@@ 3.1 Inability to Logically Inference and Robustly Generalize
When a fact is updated, it is crucial not only to revise the specific piece of knowledge but also to evaluate the impact on the related reasoning chain. Recently the term portability has been proposed in (Yao et al., 2023b) to evaluate the consequences after an edit, and further assess the robustness of generalization. In their study, they introduce three metrics to evaluate portability: Subject Replace (checking if synonyms of the subject are edited), Reversed Relation (checking if the reversed relation of the target is edited), and One Hop (assessing if modified knowledge is usable for further derivation). Similarly, RippleEdits benchmark as well as corresponding Logical Generalization and Compositionality metrics are proposed to examine whether edited knowledge can be inferred in composite relations of facts Cohen et al. (2023). Additionally, ReCoE benchmark is proposed to assess the propagation of updates in interconnected facts using various reasoning schemes in complex question-answering datasets Hua et al. (2024). Furthermore, MQuAKE benchmark is introduced to evaluate more complex reasoning and inference ability on multi-hop questions Zhong et al. (2023).
When editing multiple logically related facts simultaneously, models may suffer from confusion due to conflicts.  ConflictEdit benchmark is proposed to examine different editing methods on conflicted edit facts Li et al. (2024b). The different benchmarks and corresponding metrics and are arranged systematically in Table 2.
@@ 3.2 Unintended Alteration of Non-Target Knowledge
Locality is conventionally assessed using a locality dataset to evaluate edits on unrelated facts by measuring the Neighborhood Score and Neighborhood Magnitude (NS & NM; Meng et al., 2022, 2023).However, current evaluation methods do not adequately capture the post-edit effects on content beyond the locality dataset, which means the edited model could still contain unintended alterations. For example, while the location of the Louvre might be successfully changed from Paris to London, the edited model may inadvertently increase the likelihood of semantically related words (e.g. Big Ben for London) while mentioning Louvre. Some modified benchmark (CounterFact+) and corresponding metric (Neighborhood KL Divergence(Hoelscher-Obermaier et al., 2023) is then designed to disclose these previously implicit pitfalls. Another study (Yao et al., 2023a) extends this exploration to three facets of locality: Other Relations(evaluating the retention of other attributes of the updated subject), Distract Neighborhood(assessing the divergence of unrelated input when juxtaposed with edited instances), and Other Tasks (examining the influence of edits on the performance of other tasks).
Unintended edits to unrelated facts may occur because a single edit can implicitly change the predictive distribution among objects associated with the same (subject - relation) pair. After multiple consecutive edits, these alterations can accumulate and distort the stored knowledge. To evaluate this condition, the concept of Knowledge Distortion has been introduced by (Li et al., 2024b), which estimates the Jensen–Shannon divergence of the object set distribution before and after editing. This can be further extended to metrics such as the Ignore Rate, measuring how objects other than the target in the object set are neglected after editing, and the Failure Rate, which measures the proportion of instances where over half of the objects in the set are overlooked.

### v2

@@ Abstract
Knowledge editing is a rising technique for efficiently updating factual knowledge in large language models (LLMs) with minimal alteration of parameters. However, recent studies have identified side effects, such as knowledge distortion and the deterioration of general abilities, that have emerged after editing. Despite these findings, evaluating the pitfalls of knowledge editing often relies on inconsistent metrics and benchmarks, lacking a uniform standard. In response, this survey presents a comprehensive study of these side effects, providing a unified perspective on the challenges of knowledge editing in LLMs by conducting experiments with consistent metrics and benchmarks. Additionally, we review related works and outline potential research directions to address these limitations. Our survey highlights the limitations of current knowledge editing methods, emphasizing the need for a deeper understanding of the inner knowledge structures of LLMs and improved knowledge editing methods. To foster future research, we have released the complementary materials publicly [^1] https://github.com/MiuLab/EditLLM-Survey. **footnotetext: Equal contribution.
@@
Cheng-Hsun Hsueh Paul Kuo-Ming Huang Tzu-Han Lin Che-Wei Liao Hung-Chieh Fang Chao-Wei Huang Yun-Nung Chen National Taiwan University, Taipei, Taiwan {r12922059,b08902072,r12944034,r09922a25}@csie.ntu.edu.tw {b09902106,f07922069}@csie.ntu.edu.tw y.v.chen@ieee.org
@@ 1 Introduction
Recent advancements in large language models (LLMs) have significantly improved NLP applications, enabling LLMs to understand and generate language at a human-like level. However, the mechanisms of knowledge storage in LLMs remain unclear, raising concerns about the reliability of their output, particularly in applications like chatbots. To address these issues, researchers have explored various methods. Traditional methods like fine-tuning, continual learning, and retraining are computationally expensive and may degrade LLM performance. Knowledge editing has emerged as a promising alternative, offering efficient adjustments with minimal computational costs and fewer alterations (Cao et al., 2021; Dai et al., 2022; Meng et al., 2022, 2023; Dong et al., 2022; Mitchell et al., 2022a, b; Hartvigsen et al., 2023; Huang et al., 2023; Yu et al., 2024; Zheng et al., 2023; Li et al., 2023; Tan et al., 2024; Gupta et al., 2024b; Wang et al., 2024). This method allows precise LLMs refinement, enhancing their practical and reliable use in real-world applications.
Figure 1: An overview of pitfalls in current knowledge editing methods. The subsequent sections dive into three key challenges: generalization issues (Section 3.1), locality issues (Section 3.2), and deterioration of general LLM abilities (Section 3.3).
@@
Knowledge editing can be divided into two main categories: parameter-modifying and parameter-preserving. Both aim to refine LLM knowledge efficiently while avoiding the drawbacks of previous tuning methods (Yao et al., 2023). Parameter-modifying methods, including meta-learning (Cao et al., 2021; Mitchell et al., 2022a; Tan et al., 2024) and locate-and-edit techniques (Dai et al., 2022; Meng et al., 2022, 2023; Li et al., 2023; Gupta et al., 2024b), strive to update model parameters effectively. By contrast, parameter-preserving methods introduce external components, like knowledge bases (Mitchell et al., 2022b; Zhong et al., 2023) or extra model parameters (Dong et al., 2022; Huang et al., 2023; Hartvigsen et al., 2023; Yu et al., 2024) to maintain the integrity of pre-trained LLMs while updating their knowledge.
Despite the success of knowledge editing, challenges remain. Recent studies have revealed side effects that can harm the general capabilities and intrinsic structures of LLMs. We categorize these pitfalls into three main areas: (1) the inability to perform logical inference (Cohen et al., 2023; Li et al., 2024; Zhong et al., 2023; Hua et al., 2024; Yao et al., 2023), (2) the unintended modification of non-target knowledge (Cohen et al., 2023; Li et al., 2024; Yao et al., 2023; Meng et al., 2022; Hoelscher-Obermaier et al., 2023), and (3) the deterioration of general LLM abilities (Gupta et al., 2024a; Gu et al., 2024; Yang et al., 2024). Although various side effects have been identified, they are evaluated using inconsistent metrics and benchmarks in different studies, which lack a uniform standard. As a result, this survey aims to provide a comprehensive overview of the current issues in the knowledge editing paradigm and to establish a fair platform for comparing the side effects of different editing methods. Additionally, we encourage further investigation into the pitfalls and underlying knowledge structures of LLMs. A brief overview of the discussed pitfalls is shown in Figure  1.
@@
This paper is organized as follows: Section 2 introduces the definition and methods of knowledge editing. Section 3 discusses current challenges and corresponding benchmarks. In Section 4, we present experimental results evaluating different editing methods. Finally, Section 5 explores related studies and future research directions. We summarize our contributions as follows:
@@
    We systematically organize previous research and conduct experiments to benchmark the side effects of knowledge editing, providing a unified perspective on this issue.
@@
    We discuss related studies and potential directions to address existing challenges, encouraging further exploration in this field.
Figure 2: Illustration of properties that knowledge editing methods should satisfy: reliability, generalizability/portability, and locality.
@@ 2.1 Problem Definition
Knowledge editing for LLMs entails modifying the output of LLMs in response to specific edit queries, with the aim of minimizing alterations to their original behavior (Yao et al., 2023; Mazzia et al., 2023; Zhang et al., 2024a). In this section, we follow the notation from Mazzia et al. (2023).
@@
The editing process should not affect instances unrelated to the edit queries. The locality set of an edit query $(x_{e},y_{e})$ can be defined as $L(x_{e})=\{(x_{loc},y_{loc})\in\mathbb{X}\times\mathbb{Y}\ \mathrm{s.t}\ x_{loc}\notin N(x_{e},y_{e})\land f_{\theta_{0}}(x_{loc})=y_{loc}\}$. The locality, also known as specificity, of an editing method is measured by calculating the level of invariance of model output before and after the edits, which can be calculated as follows:
@@ 2.2 Current Methods
Current knowledge editing methods are categorized into parameter-modifying (Section 2.2.1) and parameter-preserving (Section 2.2.2) editing methods, each containing several strategies. An overview and illustration of current methods are included in Table 1 and Figure 3, respectively.
Category Strategy Method
MEND (Mitchell et al., 2022a)
MALMEN (Tan et al., 2024)
Locating and editing Knowledge Neuron (Dai et al., 2022)
ROME (Meng et al., 2022)
MEMIT (Meng et al., 2023)
PMET (Li et al., 2023)
EMMET (Gupta et al., 2024b)
T-Patcher (Huang et al., 2023)
GRACE (Hartvigsen et al., 2023)
MELO (Yu et al., 2024)
External memory SERAC (Mitchell et al., 2022b)
MeLLo (Zhong et al., 2023)
In-context learning IKE (Zheng et al., 2023)
Decoding DeepEdit (Wang et al., 2024)
Table 1: Overview of knowledge editing methods. The methods are categorized into two major families, parameter-modifying and parameter-preserving methods, each containing several strategies. Methods marked with $\dagger$ have the ability to process sequential edits.
@@ 2.2.1 Parameter-Modifying
Meta-learning methods train a hyper-network to predict network parameter updates. For instance, KnowledgeEditor (Cao et al., 2021) trains a deep network to predict weight updates. MEND (Mitchell et al., 2022a) decomposes the gradient matrix into two rank-one matrices and utilized a hyper-network to update these matrices, thereby accelerating the editing process. Built upon MEND, MALMEN (Tan et al., 2024) refines the process by formulating the aggregation of parameter shifts into a least-squares problem, further improving the scalability of meta-learning methods.
@@
Locate-and-edit methods identify specific knowledge locations in LLMs for consequent editing. KN (Dai et al., 2022) utilizes the proposed knowledge attribution method to pinpoint neurons expressing relational facts, allowing efficient updates or erasures without fine-tuning. ROME (Meng et al., 2022) proposes causal tracing method to identify neuron activations linked to specific knowledge. The authors demonstrate the significance of middle-layer feed-forward networks (FFNs) in factual predictions when processing the subject’s last token. Built upon the hypothesis that the FFN modules in a transformer layer can be viewed as key-value memories (Geva et al., 2021), ROME injects new knowledge into the key-value memories by deriving the closed form solution from the least-squares problem. MEMIT (Meng et al., 2023) scales up ROME by editing a set of MLPs from consecutive middle-layers via solving a normal equation. PMET (Li et al., 2023) proposes to update multi-head self-attention (MHSA) modules in addition to FFNs. EMMET (Gupta et al., 2024b) on the other hand, integrates the objectives of ROME and MEMIT into a unified preservation-memorization objective, facilitating batch-editing capabilities for both methodologies.
@@ 2.2.2 Parameter-Preserving
Some methods utilize additional parameters, such as adding new neurons or employing parameter-efficient techniques. CaliNET (Dong et al., 2022) extends the FFN modules with calibration memory slots to adjust the predicted token distribution. T-Patcher (Huang et al., 2023) adds neurons in the FFN’s last layer to rectify classification errors and incorrectly generated tokens, activating only in response to associated mistakes. GRACE (Hartvigsen et al., 2023) wraps a selected layer with an Adaptor that includes a codebook and deferral mechanism, learning to decode desired outputs while caching embeddings of error inputs. The GRACE layer stores the edits and could be updated continuously over long deployments. MELO (Yu et al., 2024) utilizes DyLoRA (Valipour et al., 2023) modules to learn edits, indexing them in an inner vector database to dynamically activate corresponding LoRA blocks during inference.
@@ 3 Challenges of Knowledge Editing
While knowledge editing methods have been extensively researched, comprehensive studies on related challenges are lacking. In this section, we discuss the pitfalls of knowledge editing from three perspectives: inability to logically infer and robustly generalize (Section 3.1), unintended alteration of non-target knowledge (Section 3.2), and deterioration of general LLM abilities (Section 3.3).
@@
Portability and Generalization RippleEdits (Cohen et al., 2023) Logical Generalization, Compositionality I, Compositionality II
ConflictEdit (Li et al., 2024) Conflict Score, Conflict Magnitude, Success Score
MQuAKE (Zhong et al., 2023) Edit-wise Success Rate, Instance-wise Accuracy, Multi-hop Accuracy
@@
ZsRE + CounterFact Yao et al. (2023) Subject-Replace, Reverse-Relation, One-Hop
Locality RippleEdits (Cohen et al., 2023) Subject Aliasing, Preservation, Relation Specificity
RoundEdit (Li et al., 2024) Success Score, Distortion ($\downarrow$), Ignore Rate ($\downarrow$), Failure Rate ($\downarrow$), Tied Fact Damage ($\downarrow$)
ZsRE + CounterFact Yao et al. (2023) Other-Attribution, Distract-Neighbor, Other-Task
CounterFact Meng et al. (2022) Locality, Neighborhood Score, Neighborhood Magnitude
@@
Table 2: Performance benchmarks and evaluation metrics addressing generalization/portability and locality issues in knowledge editing methods. Unless specifically indicated by a downward arrow, higher values signify better performance in those evaluation metrics. CounterFact benchmark is proposed by  (Meng et al., 2022), and CounterFact with mark is modified by  (Yao et al., 2023) to further examine the proposed metrics.
@@ 3.1 Inability to Logically Inference and Robustly Generalize
When a fact is updated, it is crucial not only to revise the specific piece of knowledge but also to evaluate the impact on the related reasoning chain. Recently the term portability has been proposed in (Yao et al., 2023) to evaluate whether an edited fact can be logically inferred within the knowledge chain, and to further assess the robustness of generalization. In their study, they introduce three metrics to evaluate portability: Subject Replace (checking if synonyms of the subject are edited), Reversed Relation (checking if the reversed relation of the target is edited), and One Hop (assessing if modified knowledge is usable for further derivation). Similarly, RippleEdits benchmark and its corresponding Logical Generalization and Compositionality metrics are proposed to examine whether edited knowledge can be inferred in composite relations of facts Cohen et al. (2023). Additionally, ReCoE benchmark is proposed to assess the propagation of updates in interconnected facts using various reasoning schemes in complex question-answering datasets Hua et al. (2024). Furthermore, MQuAKE benchmark is introduced to evaluate more complex reasoning and inference ability on multi-hop questions Zhong et al. (2023).
When multiple logically related facts are edited simultaneously, models may become confused by conflicts between their pre-existing knowledge and the newly edited information. ConflictEdit benchmark is thus proposed to examine different editing methods on conflicted edit facts Li et al. (2024). The different benchmarks and corresponding metrics and are arranged systematically in Table 2.
@@ 3.2 Unintended Alteration of Non-Target Knowledge
Locality is conventionally assessed using a locality dataset to evaluate the impact of edits on unrelated facts by measuring the Neighborhood Score and Neighborhood Magnitude (NS & NM; Meng et al., 2022, 2023). However, current evaluation methods do not adequately capture the post-edit effects on content beyond the locality dataset, which means the edited model could still contain unintended alterations. For example, while the location of the Louvre is successfully modified from Paris to London, the edited model might also output London in an unrelated context or increase the probability of words semantically related to London (e.g., Big Ben) when mentioning the Louvre. Some modified benchmark (CounterFact+) and corresponding metric (Neighborhood KL Divergence) (Hoelscher-Obermaier et al., 2023) is then designed to disclose these previously implicit pitfalls. Another study (Yao et al., 2023) extends this exploration to three facets of locality: Other Relations (evaluating the retention of other attributes of the updated subject), Distract Neighborhood (assessing whether model will be swayed by edited cases when they are concatenated before unrelated inputs), and Other Tasks (examining the influence of edits on the performance of other tasks).
Unintended edits to unrelated facts may occur because a single edit can implicitly change the predictive distribution among objects associated with the same (subject - relation) pair. After multiple consecutive edits, these alterations can accumulate and distort the stored knowledge. To evaluate this condition, the concept of Knowledge Distortion has been introduced by Li et al. (2024), which estimates the Jensen–Shannon divergence of the object set distribution before and after editing. This can be further extended to metrics such as the Ignore Rate, measuring how objects other than the target in the object set are neglected after editing, and the Failure Rate, which measures the proportion of instances where over half of the objects in the set are overlooked.